OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14 Answer -2 2 1 -1

myininaya (myininaya):

multiply second equation by -1 and then add the equations together this will eliminate x but you can solve for y and then use that y to plug into one equations given to find x

OpenStudy (anonymous):

so its .... 1?

myininaya (myininaya):

\[4x+3y=10\] \[-(4x+5y=14)\] --------------- 0-2y=-4 -2y=-4 y=2 => 4x+3(2)=10 4x+6=10 4x=10-6 4x=4 x=1

myininaya (myininaya):

yes